OpenStudy (anonymous):

What is the pattern in 2-,1,6,13,22

OpenStudy (anonymous):

n+2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

-2+3=1 1+5=6 6+7=13 13+9=22